OpenStudy (opcode):

Differentiate the function \(f(x)=x^3\) in the point a. Use the definition of the derivative for this question. I know that the definition of the derivative is: \[f'(x) = \lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\] The function \(f(x)=x^3\) Now to find the derivative... I'm just not sure how to obtain the derivative do to the function has an exponent. Here's my attempt: \[f'(x) = \lim_{h\to 0}\frac{x^3(x+x^3+x^3)-x^3}{h}\] Then I got: \[f'(x) = \lim_{h\to 0}\frac{x^3(x + 2 x^3)-x^3}{2x^3}\] Simplified: \[f'(x) = \lim_{h\to 0}= x\] I know the answer is wrong, I just cannot figure out what I did wrong.

OpenStudy (anonymous):

nonononoonnoo *joker voice*

OpenStudy (anonymous):

take the definition of derivative and put it IN YOUR FUNCTION

OpenStudy (anonymous):

\[f \prime (x) = \lim_{h \rightarrow 0} \frac{ (x+h)^3 - (x)^3}{ h }\]

OpenStudy (anonymous):

By the power rule, the answer is \(3x^2\). However, the definition: \[lim_{h\to 0}\frac{f(x+h)-f(x)}{h} \\ lim_{h\to 0}\frac{(x+h)^3-x^3}{h} \\ lim_{h\to 0}\frac{h^3+3 h^2 x+3 h x^2+x^3-x^3}{h} \\ lim_{h\to 0}\frac{h^3+3 h^2 x+3 h x^2}{h} \\ lim_{h\to 0}h^2+3 h x+3 x^2 \\ 3x^2\]

OpenStudy (anonymous):

By the way, if you'd like to expand a binomial easily, use the binomial theorem.

OpenStudy (opcode):

Oh. I see what I did wrong now *facedesk*. \[f'(x) = \lim_{h\to0}{\frac{f(x+h) - f(x)}{h}}\] \[=\lim_{h\to0}\frac{(x+h-x)((x+h)^2+(x+h)\cdot(x)+x^2)}{h}\] \[=\lim_{h\to0}((x+h)^2+(x+h)\cdot(x)+x^2)\] \[=(x)^2+(x)\cdot(x)+x^2\] \[3x^2\]